# Routine haematological markers can predict and discriminate health status and biological age even from noisy sources

### v9

@@ Abstract
For more than two decades, advances in personalised medicine and precision healthcare have largely been based on genomics and other omics data. These strategies aim to tailor interventions to individual patient profiles, promising greater treatment efficacy and more efficient allocation of healthcare resources. Here, we show that widely collected common haematologic markers can reliably predict and discriminate individual chronological age and health status from even noisy sources. Our analysis includes synthetic and real retrospective patient data, including medically relevant and extreme cases, and draws on more than 100 000 complete blood count records over 13 years from the United States Centers for Disease Control and Prevention’s National Health and Nutrition Examination Survey (CDC NHANES). We combine fully explainable risk assessment scores with machine and deep learning techniques to focus on clinically significant patterns and characteristics without functioning purely as a “black-box model allowing interpretation and control. We validated the results with the UK Biobank, a larger cohort independent of the CDC NHANES and with very different collection techniques, the former a survey and the second a longitudinal study. Unlike current biological ageing indicators, this approach may offer rapid, and scalable implementations of personalised, precision and predictive approaches to healthcare and medicine without or before requiring other specialised, uncommon or costly tests.
Keywords: Predictive blood markers, precision hematology, deep medicine, adaptive health learning, precision healthcare, predictive medicine, digital blood twin, risk stratification index, immune age, CDC NHANES, UK Biobank.
@@ 1 Introduction
In a recent retrospective study, Complete Blood Count tests of 12 000 adults over 20 years showed that individuals are stable over time while being healthy and have unique characteristics that effectively act like health fingerprints, specific patterns or personalised signatures that when deviate from their personal baselines are indicative of health and even mortality [32]. These personalised benchmarks can significantly improve early disease detection by providing a better and more precise picture of individual health. This had already been previously suggested from more specialised immunological tests [26, 27].
Here, we investigate whether these findings can be leveraged for risk assessment by identifying clinically relevant immune patterns. We further demonstrate that the same test allows robust detection of patterns predictive of age and health status, even in large, retrospective, and noisy datasets.
This study presents two main clinical goals. First, we evaluate whether immune scores derived from routine blood tests can predict and triage (i.e., assessment of priority for healthcare services) health status using a translatable and explainable approach. Second, we assess whether immune-derived age diverges from chronological age in a clinically meaningful way. The results are validated across the US NHANES and UK Biobank cohorts to assess generalisability.
@@ 1.1 Biological Ageing
In contrast, the most popular medical test, a blood test, the Complete Blood Count (CBC) or Full Blood Count, has been found to often contain enough data to provide key clinically information [19, 39, 32], and even specific disease signatures [42]. Although it is a popular low-cost laboratory test that is almost universally used, its risk predictive information content is underused [44, 32], often in favour of more sophisticated omics tests that are also expensive and are not currently scalable.
@@ 2 Methods
Thus, the numerical vector value integrates all the analyte’s spatial dimensions. We call the associated vector space for blood cell analytes the multidimensional immune space, or the immune space for short.
2.1 Numerical index
@@ 2.2 NHANES Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table S2 (see Sup. Inf.) highlights sex-specific differences in immune analyte values, for reference.
2.3 Immune Age Estimation
@@
Here, we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, provides insight into immune age compared to chronological age. Chronological age simply indicates the age of the individual, determined by the length of time someone has lived or has lived since their birth. This term should not be confused with the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
2.4 Machine Learning Methodology
2.5 Further Validation of the Indexes with the UK Biobank database
The UK Biobank is a long-term prospective nation-wide study that houses the deidentified biological samples and health-related data of half a million people. Between 2006 and 2010 volunteers aged 40 to 69 years were recruited from all over Great Britain and consented to share their health data and to be followed for at least 30 years thereafter; with the objective of allowing scientific discoveries to be made in the prevention, diagnosis, and treatment of disease [20, 22]. We tested the NIS index (which is the NCCIS index without the warning colouring component, see Supp. Inf.), to separate large disease categories in the UK Biobank. The database does not contain a clean healthy control group, meaning that all participants had potentially some reported health problem, making our separation results more relevant given that we had to create a Rest group different from the large disease categories identified as a noisy control group.
@@ 3 Results
To assess group differences in immune scores, we first conducted Shapiro–Wilk tests to evaluate normality. Due to violations of normality assumptions and the ordinal nature of the NCCIS score, we employed the non-parametric Kruskal–Wallis test to assess overall group differences. All comparisons were statistically significant (p << 0.0001). To evaluate differences between specific group pairs, we used Mann–Whitney U tests with Bonferroni correction for multiple comparisons, retaining significance across all contrasts.
@@
Healthy vs Unhealthy 7007 vs 6074 P << 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P << 0.0001
Healthy vs Anaemia 7007 vs 980 P << 0.0001
Healthy vs HIV 7007 vs 62 P << 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P << 0.0001
Healthy vs Leukaemia 7007 vs 18 P << 0.0001
Healthy vs Blood Cancer 7007 vs 8 P << 0.0001
Table 1: Differences in NIS mean values. Statistical comparisons were performed using the Kruskal–Wallis test, followed by Mann–Whitney U tests with Bonferroni correction for pairwise group comparisons.
@@
Healthy vs Unhealthy 7007 vs 6074 P << 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P << 0.0001
Healthy vs Anaemia 7007 vs 980 P << 0.0001
Healthy vs HIV 7007 vs 62 P << 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P << 0.0001
Healthy vs Leukaemia 7007 vs 18 P << 0.0001
Healthy vs Blood Cancer 7007 vs 8 P << 0.0001
Table 2: Differences in NCCIS mean values. Kruskal–Wallis tests revealed significant group differences (p << 0.0001), followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons. Effect sizes are discussed in Table 3.
@@
IMMUNE AGE (by Health) 2.28 P << 0.0001
index 1.36 P << 0.0001
IMMUNE AGE (by Sex) 0.11 P << 0.0001
Table 3: Cohen’s d effect sizes of combined blood parameters in the NIS immune index and Immune Age, comparing healthy and unhealthy groups. While Cohen’s d is robust to normality violations, we additionally report p-values using Welch’s t-test. Results demonstrate strong practical and statistical significance of the immune index.
The Immune Age by Health Status shows a strong Cohen’s d (2.28) 3, indicating that it is a strong and robust measure to differentiate health status and predict health risks based on blood parameters with precision. In contrast, the weaker effect size by sex (0.11) suggests that sex is unlikely to be a confounding factor, reinforcing the precision of the Immune Age prediction based on health status.
Figure 1: Distribution of the Normalised Colour-Coded Immune index (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (S1 (see Sup. Inf.) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. HIV is an example of chronic infection while all the diseases can fall in the category of immune and blood-related conditions.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patient groups when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 1, it is shown that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We therefore conclude that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS. Separation of groups of self-reported data suggests that they could be separated from those that were not actually healthy at the time of testing. In contrast, those who self-reported not being healthy were indeed clinically not healthy according to their results and the index. The number of individuals in each cohort is shown in table 2. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS.
@@ 3.1 Adaptive Index Behaviour Over Time
We also hypothesised that as the patient aged, the immune index would vary over time. In particular, we expected to see an upward trajectory in direct relation to age.
@@ 3.2 Discounting Confounding Factors from Age
Table 6: Statistics of each Complete Blood Count index as a predictor or discriminant of health and age.
Counts Healthy Unhealthy Age
BASO 0.01 0.02 0.00
EOS 0.00 0.00 0.00
HCT 0.00 0.01 0.07
HGB 0.12 0.14 0.06
LYMP 0.04 0.50 0.09
MCH 0.08 0.00 0.15
MCHC 0.00 0.00 0.00
MCV 0.27 0.34 0.22
MONO 0.01 0.00 0.01
NEUT 0.01 0.01 0.02
RBC 0.03 0.00 0.00
PLT 0.09 0.33 0.10
@@
Figure 5: Age Prediction Error Rates between actual patient (biological) age and Immune-index predicted age from NHANES dataset by RBC counts. The standard error analysis shown includes the prediction error rates in years (as box plots) across different health and sex groups, computed by assessing the variability in age prediction errors using interquartile ranges (IQR). The health status was determined by the best blood analyte correlate in this case as previously discussed– the RBC counts– to separate healthy from unhealthy groups by sex.
@@ 3.3 Machine Learning Guiding index
As shown in Table  10 for the different AI models, the NIS immune index, a single quantitative measure (scalar value), also showed significant classification performance, particularly when coupled with computational AI-driven predictive power using ANN networks. While other basic ML classifiers did not perform as well, the ANN paired with the NIS immune index achieved an optimal predictive accuracy of 0.987, a high specificity of 0.999, and an AUC index of 0.91, comparable to the individual analytes themselves. However, the sensitivity was poor compared to the analytes. A low sensitivity in AI classification implies it is not able to correctly classify true positives. In contrast, the basic machine learning classifiers, such as RandomForest (RF), SVM, and KNN, did not perform as well as the ANN models with the NIS immune index, achieving lower AUC scores with 10-fold cross-validation, despite their high classification accuracy. While the feedforward ANN (MLP) detects local features through its dense layers, the Transformer-ANN captures long-range dependencies using self-attention mechanisms.
Table 8 provides the highest features importance metrics (i.e., mean Gini entropy decrease) for different blood analytes using the RF algorithm. Table 9 shows the performance of all blood analytes collectively, and they perform suboptimally against ANN performance on NIS Immune scores shown in Table  10. The introduced index captures in a single numerical value the power of this and other analytes that may be of interest for other patient profiles, also giving the clinician the opportunity and option to add or remove analytes or assign different weights to different analytes. Table 11 shows that both Transformer-ANN and ANN models demonstrate superior performance with an AUC of 0.91 to predict outcomes based on the NIS index, significantly outperforming traditional models such as KNN, RF and SVM. However, their sensitivity lacks robustness and requires further optimisation through threshold adjustments, enhanced feature selection, or targeted optimisation techniques. Future studies with larger validation datasets can include bootstrapping methods and dropout layers in the ANNs to help prevent overfitting.

### v10

@@ Abstract
For more than two decades, advances in personalised medicine and precision healthcare have largely been based on genomics and other omics data. These strategies aim to tailor interventions to individual patient profiles, promising greater treatment efficacy and more efficient allocation of healthcare resources. Here, we show that widely collected common haematological markers can reliably predict and discriminate individual chronological age and health status from even noisy sources. Our analysis includes synthetic and real retrospective patient data, including medically relevant and extreme cases. We combine fully explainable risk assessment scores with machine and deep learning techniques to focus on clinically significant patterns and characteristics without functioning purely as a “black box” model allowing interpretation and control. We validated the results with the CDC NHANES database (100 000 participants) and the UK Biobank (500 000 participants), the former a noisy survey and the second a noisy longitudinal study. Unlike current biological ageing indicators, this approach may offer rapid, and scalable implementations of personalised, precision and predictive approaches to healthcare and medicine without or before requiring other specialised, uncommon or costly tests.
Keywords: Predictive blood markers, precision haematology, deep medicine, adaptive health learning, precision healthcare, predictive medicine, digital blood twin, risk stratification index, immune age.
@@ 1 Introduction
In a recent retrospective study, Complete Blood Count tests of 12 000 adults over 20 years showed that individuals are stable over time while being healthy and have unique characteristics that effectively act like health fingerprints, specific patterns or personalised signatures that when deviate from their personal baselines are indicative of health and even mortality [33]. These personalised benchmarks can significantly improve early disease detection by providing a better and more precise picture of individual health. This had already been previously suggested from more specialised immunological tests [27, 28]. Unlike the study [33], the present work focusses on defining a minimal, interpretable index or risk assessment score grounded in routine haematological and circulatory biomarkers. Taking advantage of population-normalised Complete Blood Count (CBC) parameters, our approach achieves discriminatory performance not explored in [33] while remaining computationally efficient, reproducible and transparent for clinical purposes. The index therefore is the proactive natural consequence of what was reported in  [33] regarding personal setpoints and clinically relevant deviations even within normal population reference values.
Here, we investigate whether these findings can be leveraged for risk assessment by identifying clinically relevant immune patterns. We further demonstrate that the same test allows robust detection of patterns predictive of age and health status, even in large, retrospective, and noisy datasets. In this paper [1], it was shown that 60 blood markers were predictive of biological age. While our paper precedes theirs as posted online, in this one we show that only a small subset of those markers are actually necessary for predictive purposes and that ad hoc artificial weighting of markers is not necessary.
This contribution has two main goals. First, we evaluate whether immune scores derived from routine blood tests can predict and triage (i.e., assessment of priority for healthcare services) health status using a translatable and explainable approach. Second, we assess whether immune-derived age diverges from chronological age in a clinically meaningful way. The results are validated across the US NHANES and UK Biobank cohorts to assess generalisability.
@@ 1.1 Biological Ageing
In contrast, the most popular medical test, a blood test, the CBC or Full Blood Count, has been found to often contain enough data to provide key clinically information [20, 40, 33], and even specific disease signatures [43]. Although it is a popular low-cost laboratory test that is almost universally used, its risk predictive information content is underused [45, 33], often in favour of more sophisticated omics tests that are also expensive and are not currently scalable.
@@ 2 Methods
Thus, the numerical vector value integrates all the analyte’s spatial dimensions. We call the associated vector space for blood cell analytes the multidimensional immune space, or immune space for short.
2.1 Numerical index and Age Prediction
The following abbreviations are used throughout the paper: NIS (Numerical Immune Score), CCIS (Colour-Coded Immune Score), NCCIS (Normalised Colour-Coded Immune Score), and Immune Index (collective term encompassing all variants). In brief, the NIS represents the raw numerical formulation based solely on analyte deviations; the CCIS extends this by implementing a categorical colour-coding system for visually assisted interpretability; and the NCCIS normalises these values across population-level reference intervals, improving cross-cohort comparability. The Immune Index denotes the general framework comprehensive of all three forms.(See Table 1)
Abbreviation Description
NIS (Numerical Immune Score) Quantitative index derived from haematological analyte deviations; represents raw numerical immune deviation without categorical scaling.
CCIS (Colour-Coded Immune Score) Categorical extension of NIS using colour-coded bins to improve interpretability and visual comparison between individuals.
NCCIS (Normalised Colour-Coded Immune Score) Normalised version of CCIS adjusted to population reference intervals (NHANES, NHS-UK), allowing generalisation and comparison across diverse datasets and laboratories.
Immune Index Collective term referring to the overall framework encompassing NIS, CCIS, and NCCIS for generalisable immune health quantification.
Table 1: Summary of immune-related indices and their definitions. Each variant of the immune index represents a step toward increased interpretability and cross-dataset comparability, forming a hierarchical structure from raw numerical scoring (NIS) using the CBC analytes to population-normalised colour-coded indices (NCCIS).
@@ 2.1 Numerical index and Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. Zero values were excluded only when they represented non-physiological or missing data artifacts (e.g., unrecorded analytes). True physiological zeros, where relevant, were retained. However, for the quantitative analysis, sex was replaced by 1 for males and 2 to denote females. Table S2 (see Sup. Inf.) highlights sex-specific differences in immune analyte values, for reference. Because the NHANES and UK Biobank datasets compile data from multiple laboratories, we standardized all analyte values using the corresponding national population reference intervals (NHANES and NHS-UK) prior to computing the indices. This population-level normalization mitigates inter-laboratory cut-off variability and supports cross-site comparability.
@@
Here, we present a method where a single blood test, as in a Complete Blood Count (CBC) of 13 blood analytes, provides insight into immune age compared to chronological age. Chronological age simply indicates the age of the individual, determined by the length of time someone has lived or has lived since their birth. This term should not be confused with the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
The UK Biobank is a long-term prospective nationwide study that houses the deidentified biological samples and health-related data of half a million people. Between 2006 and 2010 volunteers aged 40 to 69 years were recruited from all over Great Britain and consented to share their health data and to be followed for at least 30 years thereafter; with the objective of allowing scientific discoveries to be made in the prevention, diagnosis, and treatment of disease [21, 23]. We tested the NIS index (which is the NCCIS index without the warning colouring component, see Supp. Inf.), to separate large disease categories in the UK Biobank. The database does not contain a clean healthy control group, meaning that all participants had potentially some reported health problem, making our separation results more relevant given that we had to create a Rest group different from the large disease categories identified as a noisy control group.
@@ 3 Results
To assess group differences in immune scores, we first conducted a graphical analysis of the residuals followed by Shapiro–Wilk tests to evaluate normality. Due to violations of normality assumptions and the ordinal nature of the NCCIS score, we employed the non-parametric Kruskal–Wallis test to assess overall group differences. All comparisons were statistically significant (p << 0.0001). To evaluate differences between specific group pairs, we used Mann–Whitney U tests with Bonferroni correction for multiple comparisons, retaining significance across all contrasts.
@@
Healthy vs Unhealthy 7007 vs 6074 <0.0001<0.0001
Self-reported Healthy vs Unhealthy 11976 vs 6074 <0.0001<0.0001
Healthy vs Anaemia 7007 vs 980 <0.0001<0.0001
Healthy vs HIV 7007 vs 62 <0.0001<0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 <0.0001<0.0001
Healthy vs Leukaemia 7007 vs 18 <0.0001<0.0001
Healthy vs Blood Cancer 7007 vs 8 <0.0001<0.0001
Table 2: Group-wise comparison of NIS mean values. Kruskal–Wallis tests revealed significant differences across all groups, followed by Mann–Whitney U tests with Bonferroni correction for pairwise comparisons. Values illustrate the robustness of the NIS in differentiating between healthy and disease cohorts.
@@
Healthy vs Unhealthy 7007 vs 6074 <0.0001<0.0001
Self-reported Healthy vs Unhealthy 11976 vs 6074 <0.0001<0.0001
Healthy vs Anaemia 7007 vs 980 <0.0001<0.0001
Healthy vs HIV 7007 vs 62 <0.0001<0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 <0.0001<0.0001
Healthy vs Leukaemia 7007 vs 18 <0.0001<0.0001
Healthy vs Blood Cancer 7007 vs 8 <0.0001<0.0001
Table 3: Group-wise comparison of NCCIS mean values. Kruskal–Wallis tests revealed significant differences across all cohorts, followed by Bonferroni-corrected Mann–Whitney U tests for pairwise comparisons. These results confirm the NCCIS as a robust and generalizable metric for distinguishing immune and haematological states. Effect sizes are provided in Table 4.
@@
Immune Age (by Health) 2.28 <0.0001<0.0001
NIS Index (by Health) 1.36 <0.0001<0.0001
Immune Age (by Sex) 0.11 <0.0001<0.0001
Table 4: Effect sizes of immune-related indices. Cohen’s d values quantify the magnitude of group differences in immune parameters across health and sex categories. Results demonstrate strong statistical and practical significance of the immune indices, highlighting their discriminative capacity for biological age and health status.
The Immune Age by Health Status shows a strong Cohen’s d (2.28) 4, indicating that it is a strong and robust measure to differentiate health status and predict health risks based on blood parameters with precision. In contrast, the weaker effect size by sex (0.11) suggests that sex is unlikely to be a confounding factor, reinforcing the precision of the Immune Age prediction based on health status.
Figure 1: Distribution of the Normalised Colour-Coded Immune index (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (S1 (see Sup. Inf.) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. HIV is an example of chronic infection while all the diseases can fall in the category of immune and blood-related conditions.While separation between overtly healthy and disease groups may be expected, the novelty of our proposed metric lies in its ability to provide a generalizable, quantitative, and interpretable single numerical score that scales across heterogeneous conditions and noisy datasets, offering a standardized baseline for predictive stratification.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patient groups when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 1, it is shown that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We therefore conclude that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS.Separation of groups of self-reported data suggests that they could be separated from those that were not actually healthy at the time of testing. The self-reported health information may not always correspond to objectively verified clinical status. Thus, the observation possibly reflects statistical discrepancies between self-reported and haematological patterns rather than confirmed diagnostic misclassification. Accordingly, while individuals who self-reported not being healthy generally exhibited blood profiles consistent with non-healthy patterns, this is suggested as a potential interpretation rather than clinical distinction. The number of individuals in each cohort is shown in Table 3. One possible explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS.
@@ 3.1 Adaptive Index Behaviour Over Time
We also hypothesised that as the patient ages, the immune index would vary over time. In particular, we expected to see an upward trajectory in direct relation to age.
@@ 3.2 Discounting Confounding Factors from Age
Table 7: Statistics of each Complete Blood Count (CBC) analyte as a predictor or discriminant of health and age.
Analyte Full Name Unit Healthy Unhealthy Age
BASO Basophils $10^{-9}$/L 0.01 0.02 0.00
EOS Eosinophils $10^{-9}$/L 0.00 0.00 0.00
HCT Haematocrit L/L 0.00 0.01 0.07
HGB Haemoglobin g/L 0.12 0.14 0.06
LYMP Lymphocytes $10^{-9}$/L 0.04 0.50 0.09
MCH Mean Cell Haemoglobin pg 0.08 0.00 0.15
MCHC MCH Concentration g/L 0.00 0.00 0.00
MCV Mean Cell Volume fL 0.27 0.34 0.22
MONO Monocytes $10^{-9}$/L 0.01 0.00 0.01
NEUT Neutrophils $10^{-9}$/L 0.01 0.01 0.02
RBC Red Blood Cells $10^{-12}$/L 0.03 0.00 0.00
PLT Platelets $10^{-9}$/L 0.09 0.33 0.10
@@
Figure 5: Age Prediction Error Rates between actual patient (biological) age and Immune-index predicted age from NHANES dataset by CBC counts. The standard error analysis shown includes the prediction error rates in years (as box plots) across different health and sex groups, computed by assessing the variability in age prediction errors using interquartile ranges (IQR). The correlation between predicted (immune) and chronological age, and the associated error distribution, are depicted in Fig. 5, which functionally serves as a scatter representation of these relationships.The health status was determined by the best blood analyte correlate to separate healthy from unhealthy groups by sex.
@@ 3.3 Machine Learning Guiding index
We explored how we could exploit the strong results from using the index as a risk assessment score and biological age predictor combined with traditional Machine Learning methods. We found that combining the symbolic strengths of a risk assessment index with the machine learning methods helped the algorithms to perform better while making the combined neurosymbolic result, a more interpretable approach given the explainability power of the index alone.
@@
As shown in Table  11 for the different AI models, the NIS immune index, a single quantitative measure (scalar value), also showed significant classification performance, particularly when coupled with computational AI-driven predictive power using ANN networks. While other basic ML classifiers did not perform as well, the ANN paired with the NIS immune index achieved an optimal predictive accuracy of 0.987, a high specificity of 0.999, and an AUC index of 0.91, comparable to the individual analytes themselves. However, the sensitivity was poor compared to the analytes. A low sensitivity in AI classification implies it is not able to correctly classify true positives. In contrast, the basic machine learning classifiers, such as RandomForest (RF), SVM, and KNN, did not perform as well as the ANN models with the NIS immune index, achieving lower AUC scores with 10-fold cross-validation, despite their high classification accuracy. While the feedforward ANN (MLP) detects local features through its dense layers, the Transformer-ANN captures long-range dependencies using self-attention mechanisms.
Table 9 provides the highest features importance metrics (i.e., mean Gini entropy decrease) for different blood analytes using the RF algorithm. Table 10 shows the performance of all blood analytes collectively, and they perform suboptimally against ANN performance on NIS Immune scores shown in Table  11. The introduced index captures in a single numerical value the power of this and other analytes that may be of interest for other patient profiles, also giving the clinician the opportunity and option to add or remove analytes or assign different weights to different analytes. Table 12 shows that both Transformer-ANN and ANN models demonstrate superior performance with an AUC of 0.91 to predict outcomes based on the NIS index, significantly outperforming traditional models such as KNN, RF and SVM. However, their sensitivity lacks robustness and requires further optimisation through threshold adjustments, enhanced feature selection, or targeted optimisation techniques. Future studies with larger validation datasets can include bootstrapping methods and dropout layers in the ANNs to help prevent overfitting.